okey, so my homework has this problem and i am a bit confused. it says that this girl drives from houston to austin is 170 miles. if she averages 67.5 miles per hour, how long will it take her to drive to austin?
Answer by oberobic(2304)

Remember the distance formula: d=rt
where
d= distance
r= rate (or you might think of it as speed)
t= time
.
Also remember that the units of measure will often help you figure our what goes where.
d = miles
r = miles/hr
t = hr
So, when you multiple r*t, the "hr" cancel and you end up with miles.
.
The setup is:
d = 170 miles
r = 67.5 miles/hr
t = ??
.
Plug in what you know.
170 miles = 67.5 miles/hr * t hrs
Divide both sides by 67.5 mi/hr
170/67.5 = t
t = 170/67.5
t = 2.5185182 hr
.
Done.
.